A sample of 49 eggs yields a mean weight of 1.69 ounces.
Assume sigma=0.46.
Find the margin of error in estimating μ at the 95% level of confidence.

Answer by stanbon(75887)

A sample of 49 eggs yields a mean weight of 1.69 ounces.
Assume sigma=0.46.
Find the margin of error in estimating μ at the 95% level of confidence.
-----
ME = z*s/sqrt(n) = 1.96*0.46/sqrt(49) = 0.1288
-----------
Cheers,
Stan H.
-------